How do I show that two given vectors are orthogonal
1
Expert's answer
2020-11-24T05:28:46-0500

Two vectors are orthogonal when an angle between them is "\frac{\pi}{2}" in radians

or 90 degrees. Then scalar production equals 0, because "\cos{\frac{\pi}{2}}=0."

If we have two vectors "\overrightarrow{x}(x_1,x_2)" and "\overrightarrow{y}(y_1,y_2)" in Euclidean system

then "x_1*y_1+x_2*y_2=0."